# Advantage Alignment Algorithms

### v1

@@ 4.1 Interpreting Advantage Alignment
Let the time dependent sets of gradients of log probabilities $B_{t}:=\{\nabla_{\theta_{1}}\log\pi^{1}(a_{0}|s_{0}),\ldots,\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\}$ and $C_{t}:=\{\nabla_{\theta_{2}}\log\pi^{2}(b_{0}|s_{0}),\ldots,\>\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})\}$ contain all the gradients up to time $t$. Then the opponent shaping term in LOLA’s update can be rewritten as
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}A^{2}_{t}\left(\sum_{\nabla_{a}\in B_{t}}\sum_{\nabla_{b}\in C_{t}}\nabla_{a}\nabla_{b}^{T}\right)\right]\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}A^{1}_{t}\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})\right].$ (10)
The proof can be found in Appendix A.2.
Theorem 2 (LOQA as an advantage aligment estimator).
The opponent shaping term in LOQA is equivalent to the opponent shaping term in advantage alignment up to a scalar $(1-\tilde{\pi}^{2}(b_{t}|s_{t}))$
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}(1-\tilde{\pi}^{2}(b_{t}|s_{t}))\cdot\gamma^{k}A^{1}_{t}A^{2}_{k}\nabla_{\theta^{1}}\log\pi^{1}(a_{k}|s_{k})\right]$ (11)
where $\tilde{\pi}^{2}(b_{t}|s_{t})$ approximates the opponent policy as defined in LOQA. For a proof see appendix A.4.
@@ 5.1 Iterated Prisoner’s Dilemma
We consider the full history version of IPD, where a gated recurrent unit (GRU) policy conditions on the full trajectory of observations before sampling an action. In this experiment we follow the architecture used in POLA (Zhao et al.,, 2022) (for details see appendix B.1). We also consider trajectories of length 16 with a discount factor, $\gamma$, of 0.9. As shown in figure 1, Advantage Alignment agents consistently achieve a policy that resembles tit-for-tat (Rapoport and Chammah,, 1965) empirically. Tit-for-tat consists of cooperating on the first move and then mimicking the opponent’s previous move in subsequent rounds.
Figure 1: Empirical probability of cooperation for each previous combination of actions in the Iterated Prisoner’s Dilemma. The results are averaged over 10 random seeds, the black whiskers show a one standard deviation confidence interval.
@@ 5.2 Coin Game
The Coin Game is a 3x3 grid world environment where two agents, red and blue, take turns collecting coins. During each turn, a coin of either red or blue color spawns at a random location on the grid. Agents receive a reward of +1 for collecting any coin but incur a penalty of -3 if the opponent collects a coin of their color. A Pareto-optimal strategy in the Coin Game is for each agent to collect only the coins matching their color, as this approach maximizes the total returns for both agents.
@@
Figure 2 demonstrates that Advantage Alignment agents perform similarly to LOQA agents when evaluated against a league of different policies: Advantage Alignment agents cooperate with themselves, cooperate with always cooperate (AC) and are not exploited by always defect (AD). In contrast to LOQA, Advantage Alignment allows for different flavors of agents. Figure 3 shows the behavior of different Advantage Alignment training by masking different quadrants. As expected, agents get more defective when the Cooperative (C) and Empathetic (E) quadrants are masked, and more cooperative when the Vengeful (V) and Spiteful (S) quadrants are masked.
Figure 3: League Results of the different ablations of Advantage Alignment agents by masking different quadrants in the Coin Game. The letters specify the quadrants that were masked for each variation: Cooperative (C), Empathetic (E), Vengeful (V) and Spiteful (S).
@@ 5.3 Negotiation Game
In the original Negotiation Game, two agents bargain over $n$ types of items over multiple rounds. In each round, both the quantity of items and the value each agent places on them are randomly set, but the agents only know their own values. They take turns proposing how to divide the items over a random number of turns. Agents can end the negotiation by agreeing to a proposal, and rewards are based on how well the agreement matches their private values. If they don’t reach an agreement by the final turn, neither gets a reward.
We modify the game first by making the values public, otherwise Advantage Alignment would have an unfair edge over PPO agents by using the opponent’s value function. Secondly, we do one-shot, simultaneous negotiations instead of negotiation rounds lasting multiple iterations. Third, we modify the reward function so that every negotiation yields a reward. For a given item with agent value $v_{a}$, the reward of the agent $r_{a}$ depends on the proposal of the agent $p_{a}$ and the proposal of the opponent $p_{o}$ where $p_{a},p_{o}\in[0,5]$:
$r_{a}=\frac{p_{a}\cdot v_{a}}{\max(5,p_{a}+p_{o})}$
Note that the max operation at the denominator serves to break the invariance of the game dynamics to the scale of proposals. For example, without the max operation, there would be no difference between $p_{a}=1,p_{o}=1$ and $p_{a}=5,p_{o}=5$.
@@
The social dilemma in this version of the negotiation game arises because both agents are incentivized to take as many items as possible, but by doing so, they end up with a lower return compared to the outcome they would achieve if they split the items based on their individual utilities. A Pareto-optimal strategy entails allowing the agent to take all the items that are more valuable to them, and similarly for their opponent (this constitutes the always cooperate (AC) strategy in Figure 4(a)). We experiment with a high-contrast setting where the utilities of objects for the agents are orthogonal to each other: There are two possible combinations of values in this setup: $v_{a}=5,v_{b}=1$ or $v_{a}=1,v_{b}=5$.
@@
As shown in Figure 4(a), PPO agents do not learn to solve the social dilemma. They learn the naive policy of bidding high for every item which means they get a low return against themselves. PPO agents trained with shared rewards get a high return against themselves, only to be exploited by PPO agents. They do not learn to abandon cooperation and retaliate after they are defected against. Advantage Alignment agents solved the social dilemma. They cooperate with themselves while remaining non-exploitable against Always Defect.
@@
Figure 4: (a) League Results of the Advantage Alignment agents in the Negotiation Game: Always Cooperate (AC), an agent which proposes $5$ for items which are more valuable to it and $1$ for items that are less valuable to it, Always Defect (AD), an agent that proposes $5$ regardless of the values, Advantage Alignment (AdAlign), PPO and PPO summing rewards (PPO-SR). Each number in the plot is computed by running 10 random seeds of each agent head to head with 10 seeds of another for 50 episodes of length 16 and averaging the rewards. Note that in this game, Always cooperate which is the most exploitable against the Always Defect gets an average return of $0.25$ and the least Always Defect as the least exploitable agent achieves $0.30$. While PPO and PPO-SR both fail to solve the game by having a cooperative reciprocation-based policy, Advantage Alignment largely solves it. (b) Sample trajectories of AdAlign vs. AdAlign and PPO vs. PPO in the negotiation game. The numbers show the utilities and proposals, which have been rounded to integer values. AdAlign agents defect first (red) and progressively cooperate whith each other (green) while PPO agents always defect.
@@ 6 Related Work
Opponent shaping is first introduced in LOLA Foerster et al., 2018b , as a method for controlling the learning dynamics of opponents in a game. To do so, a LOLA agent assumes the opponents are naive learners and differentiates through a one step look-ahead optimization update of the opponent. More formally, LOLA maximizes $V^{1}(\theta^{1},\theta^{2}+\Delta\theta^{2})$ where $\Delta\theta^{2}$ is a naive learning step in the direction that maximizes the opponent’s value function $V^{2}(\theta^{1},\theta^{2})$. Variations of LOLA have been introduced to have formal stability guarantees (Letcher et al.,, 2021), learn consistent update functions assuming mutual opponent shaping (Willi et al.,, 2022) and be invariant to policy parameterization (Zhao et al.,, 2022). More recent work performs opponent shaping by having an agent play against a best response approximation of their policy (Aghajohari et al., 2024a, ). LOQA (Aghajohari et al., 2024b, ), on which this work is based, performs opponent shaping by controlling the Q-values of the opponent using REINFORCE (Williams,, 1992) estimators.
@@
Another approach to finding socially beneficial equilibria in general sum games relies on modeling the problem as a meta-game, where meta-rewards correspond to the returns on the inner game, meta-states correspond to joint policies of the players, and the meta-actions are updates to these policies. Al-Shedivat et al., (2018) introduce a continuous adaptation framework for multi-task learning that uses meta-learning to deal with non-stationary environments. MFOS (Lu et al.,, 2022) uses model-free optimization methods like PPO and genetic algorithms to optimize the meta-value of the meta-game. More recently Meta-Value Learning (Cooijmans et al.,, 2023) parameterizes the meta-value as a neural network and applies Q-learning to capture the future effects of changes to the inner policies.
@@
The Negotiation Game, introduced by DeVault et al., (2015); Lewis et al., (2017) and subsequently refined by Cao et al., (2018), has proven to be a significant benchmark for studying general-sum games. it integrates elements of strategy and social dilemmas, necessitating that agents balance cooperation and competition to optimize their outcomes. Noukhovitch et al., (2021) analyze this complex benchmark, underscoring its importance in the field. Future investigations will turn towards an even more sophisticated simulation proposed by Zhang et al., (2022), which involves negotiations among countries and regions with diverse resource distributions and preferences in addressing climate change. This advanced model aims to provide deeper insights into negotiation strategies and outcomes in intricate, real-world scenarios, potentially informing policy and decision-making processes in global climate negotiations.
@@ 7 Conclusion
We introduced Advantage Alignment, a family of algorithms designed to shape the behavior of reinforcement learning opponents in social dilemmas. Advantage Alignment agents are able to find cooperative and non-exploitable policies, by aligning the advantages of different players and increasing the log probability of actions proportionally to this alignment.
We have also shown that LOLA and LOQA implicitly perform Advantage Alignment through different mechanisms, thus unifying and elucidating existing opponent shaping techniques. Our approach eliminates the need for imagined parameter updates or complex stochastic gradient estimation techniques, which are common in algorithms like LOLA, POLA, and LOQA. In the context of the Negotiation Game, a complex social dilemma scenario, we demonstrated the effectiveness of Advantage Alignment. Our algorithm achieved state-of-the-art results in the continuous action version of the Negotiation Game, showcasing the potential of Advantage Alignment agents to handle real-world applications such as formulating climate negotiation strategies.

### v3

@@ 4.2 Analyzing Advantage Alignment
Given a two-player game where players 1 and 2 have respective policies $\pi^{1}(a|s)$ and $\pi^{2}(b|s)$, where each policy is parametrised such that the set of gradients $\nabla_{\theta_{2}}\log\pi^{2}(a|s)$ for all pairs $(a,s)$ form an orthonormal basis, the LOLA update for the first player correspond to a reinforce update with the following opponent shaping term
$\beta\cdot\mathbb{E}_{\tau\sim\Pr_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\sum_{k=t}^{\infty}d_{\gamma,{k-t}}\gamma^{k-t}A^{1}_{k}A^{2}_{k-t}\right)\nabla_{\theta^{1}}\log\pi^{1}(a_{t}|s_{t})\right],$ (11)
where $A^{i}_{k}:=A^{i}(s_{k},a_{k},b_{k})$ and $d_{\gamma,k}$ is the occupancy measure of the tuple $(a_{k},b_{k},s_{k})$ and $\beta$ is the step size of the naive learner. See appendix A.2 for a proof.
Theorem 2 (LOQA as an advantage alignment estimator).
Under Assumption 2, the opponent shaping term in LOQA is equivalent to the opponent shaping term in Equation 8 up to $(1-\tilde{\pi}^{2}(b_{k}|s_{k}))$
$\beta\cdot\mathbb{E}_{\tau\sim\Pr_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t+1}\left(\sum_{k<t}^{\infty}\gamma^{t-k}(1-\tilde{\pi}^{2}(b_{k}|s_{k}))A^{1}_{k}\right)A^{2}_{t}\nabla_{\theta^{1}}\log\pi^{1}(a_{t}|s_{t})\right],$ (12)
$\tilde{\pi}^{2}(b_{t}|s_{t})$ approximates the opponent policy as defined in LOQA. For a proof see appendix A.5.
Having established the connection between existing opponent shaping algorithms and Advantage Alignment, we now focus on analyzing the theoretical properties of Advantage Alignment itself. We investigate the impact of the Advantage Alignment term on Nash equilibria. Theorem 3 demonstrates that Advantage Alignment preserves Nash equilibria, ensuring that if agents are already playing equilibrium strategies, the Advantage Alignment updates will not cause the policy gradient to deviate from them locally.
Theorem 3 (Advantage Alignment preserves Nash equilibria).
Advantage Alignment preserves Nash equilibria. That is, if a joint policy $(\pi_{1}^{*},\pi_{2}^{*})$ constitutes a Nash equilibrium, then applying the Advantage Alignment formula will not change the policy, as the gradient contribution of the advantage alignment term is zero. The proof can be found in Appendix A.8.
@@ 5 Experiments
We follow the evaluation protocol of LOQA (Aghajohari et al., 2024b, ), where the fixed policy that is generated by the algorithm is evaluated zero-shot against a distribution of policies.
@@ 5.1 Iterated Prisoner’s Dilemma
We consider the full history version of IPD, where a gated recurrent unit (GRU) policy conditions on the full trajectory of observations before sampling an action. In this experiment we follow the architecture used in POLA (Zhao et al.,, 2022) (for details see appendix B.1). We also consider trajectories of length 16 with a discount factor, $\gamma$, of 0.9. As shown in figure 1(b), Advantage Alignment agents consistently achieve a policy that resembles tit-for-tat (Rapoport and Chammah,, 1965) empirically. Tit-for-tat consists of cooperating on the first move and then mimicking the opponent’s previous move in subsequent rounds.
@@ 5.2 Coin Game
The Coin Game is a 3x3 grid world environment where two agents, red and blue, take turns collecting coins. During each turn, a coin of either red or blue color spawns at a random location on the grid. Agents receive a reward of +1 for collecting any coin but incur a penalty of -3 if the opponent collects a coin of their color. A Pareto-optimal strategy in the Coin Game is for each agent to collect only the coins matching their color, as this approach maximizes the total returns for both agents. Figure 2 demonstrates that Advantage Alignment agents perform similarly to LOQA agents when evaluated against a league of different policies: Advantage Alignment agents cooperate with themselves, cooperate with Always Cooperate (AC) and are not exploited by Always Defect (AD).
@@ 5.3 Negotiation Game
In the original Negotiation Game, two agents bargain over $n$ types of items over multiple rounds. In each round, both the quantity of items and the value each agent places on them are randomly set, but the agents only know their own values. They take turns proposing how to divide the items over a random number of turns. Agents can end the negotiation by agreeing to a proposal, and rewards are based on how well the agreement matches their private values. If they don’t reach an agreement by the final turn, neither gets a reward. We modify the game first by making the values public, otherwise Advantage Alignment would have an unfair edge over PPO agents by using the opponent’s value function. Secondly, we do one-shot, simultaneous negotiations instead of negotiation rounds lasting multiple iterations. Third, we modify the reward function so that every negotiation yields a reward. For a given item with agent value $v_{a}$, the reward of the agent $r_{a}$ depends on the proposal of the agent $p_{a}$ and the proposal of the opponent $p_{o}$ where $p_{a},p_{o}\in[0,5]$: $r_{a}=p_{a}\cdot v_{a}/\max(5,p_{a}+p_{o}).$
@@
Note that the max operation at the denominator serves to break the invariance of the game dynamics to the scale of proposals. For example, without the max operation, there would be no difference between $p_{a}=1,p_{o}=1$ and $p_{a}=5,p_{o}=5$. The social dilemma in this version of the negotiation game arises because both agents are incentivized to take as many items as possible, but by doing so, they end up with a lower return compared to the outcome they would achieve if they split the items based on their individual utilities. A Pareto-optimal strategy entails allowing the agent to take all the items that are more valuable to them, and similarly for their opponent (this constitutes the Always Cooperate (AC) strategy in Figure 3(a)). We experiment with a high-contrast setting where the utilities of objects for the agents are orthogonal to each other: There are two possible combinations of values in this setup: $v_{a}=5,v_{b}=1$ or $v_{a}=1,v_{b}=5$.
@@
As shown in Figure 3(a), PPO agents do not learn to solve the social dilemma. They learn the naive policy of bidding high for every item which means they get a low return against themselves. PPO agents trained with shared rewards get a high return against themselves, only to be exploited by PPO agents. They do not learn to abandon cooperation and retaliate after they are defected against. Advantage Alignment agents solves the social dilemma. They cooperate with themselves while remaining non-exploitable against Always Defect.
@@
Figure 3: (a) League Results of the Advantage Alignment agents in the Negotiation Game: Always Cooperate (AC), an agent which proposes $5$ for items which are more valuable to it and $1$ for items that are less valuable to it, Always Defect (AD), an agent that proposes $5$ regardless of the values, Advantage Alignment (AdAlign), PPO and PPO summing rewards (PPO-SR). Each number in the plot is computed by running 10 random seeds of each agent head to head with 10 seeds of another for 50 episodes of length 16 and averaging the rewards. Note that against Always Defect, Always Cooperate gets an average return of $0.25$ while Always Defect gets $0.30$. (b) Sample trajectories of AdAlign vs. AdAlign and PPO vs. PPO in the negotiation game. The numbers show the utilities and proposals, which have been rounded to integer values. AdAlign agents defect first (red) and progressively cooperate with each other (blue) while PPO agents Always Defect.
5.4 Melting Pot’s Commons Harvest Open
Figure 4: Comparison of different reinforcement learning algorithms in Melting Pot’s 2.0. Commons Harvest Open. The score is the focal return per capita, min-max normalized between a random agent and an exploiter baseline (ACB agent with an LSTM policy/value network) trained for $10^{9}$ steps. Following the protocol of the Melting Pot contest, we select the best agent out of 10 seeds and evaluate it 100 times.
In Commons Harvest Open (Agapiou et al.,, 2023), a group of 7 agents interact in a environment in which there is 6 bushes with different amounts of apples. Agents receive a reward of 1 for any apple consumed. Consumed apples regrow with a probability dependent on the number of apples in their $L_{2}$ neighborhood; specifically, if there are no apples nearby, consumed apples do not regrow. This mechanism creates a tension between agents: they must exercise restraint to prevent extinction while also feeling compelled to consume quickly out of fear that others may over-harvest.
There are a number of complications that make the Melting Pot environments particularly challenging. First, the environments are partially-observable: agents can only see a local window around themselves. Second, the partial observations are in the form of high-dimensional raw pixel data. Third, these environments often involve multiple agents—seven in the case of Commons Harvest Open—which increases the complexity of interactions and coordination. Therefore, agents need to remember past interactions with other agents to infer their motives and policies. All these factors, combined with the inherent social dilemma reward structure of the game, make finding policies that are optimal with respect to social and individual objectives a non-trivial task.
We train a GTrXL transformer (Parisotto et al.,, 2019) for 34k steps, with context length of 30, and compare the normalized focal return per capita of our agents against the baselines in Melting Pot 2.0: Advantage-critic baseline (acb) (Espeholt et al.,, 2018), V-MPO (vmpo) (Song et al.,, 2019), options as responses (opre) (Vezhnevets et al.,, 2020), and prosocial versions of opre (opre_p) and acb (acb_p) that encourage cooperation. We also compare to our own implementations of PPO (ppo) and PPO with summed rewards (ppo_p). Figure 4 shows that our best advantage alignment agent achieves on average $1.63$ normalized per capita focal return in the Commons Harvest evaluation scenarios, significantly outperforming all baselines (see Appendix B.4). Figure 5, qualitatively shows the reason why Proximal Advantage Alignment outperforms PPO and PPO with summed rewards on one of the evaluation scenarios.
Figure 5: Frames of evaluation trajectories for different algorithms. Qualitatively, we demonstrate that Proximal Advantage Alignment (adalign) also outperforms naive PPO (ppo) and PPO with summed rewards. The evaluation trajectories show how adalign agents are able to maintain a bigger number of apple bushes from extinction (2) for a longer time that either ppo or ppo_p. Note that in the Commons Harvest evaluation two exploiter agents, green and yellow, play against a focal population of 5 copies of the evaluated algorithm.
@@ 6 Related Work
Opponent shaping was first introduced in LOLA Foerster et al., 2018b , as a method for controlling the learning dynamics of opponents in a game. A LOLA agent assumes the opponents are naive learners and differentiates through a one step look-ahead optimization update of the opponent. More formally, LOLA maximizes $V^{1}(\theta^{1},\theta^{2}+\Delta\theta^{2})$ where $\Delta\theta^{2}$ is a naive learning step in the direction that maximizes the opponent’s value function $V^{2}(\theta^{1},\theta^{2})$. Variations of LOLA have been introduced to have formal stability guarantees (Letcher et al.,, 2021), learn consistent update functions assuming mutual opponent shaping (Willi et al.,, 2022) and be invariant to policy parameterization (Zhao et al.,, 2022). More recent work performs opponent shaping by having an agent play against a best response approximation of their policy (Aghajohari et al., 2024a, ). LOQA (Aghajohari et al., 2024b, ), on which this work is based, performs opponent shaping by controlling the Q-values of the opponent using REINFORCE (Williams,, 1992) estimators.
@@
Another approach to finding socially beneficial equilibria in general sum games relies on modeling the problem as a meta-game, where meta-rewards correspond to the returns on the inner game, meta-states correspond to joint policies of the players, and the meta-actions are updates to these policies. Al-Shedivat et al., (2018) introduce a continuous adaptation framework for multi-task learning that uses meta-learning to deal with non-stationary environments. MFOS (Lu et al.,, 2022) uses model-free optimization methods like PPO and genetic algorithms to optimize the meta-value of the meta-game. More recently Meta-Value Learning (Cooijmans et al.,, 2023) parameterizes the meta-value as a neural network and applies Q-learning to capture the future effects of changes to the inner policies. Shaper (Khan et al.,, 2024), scales opponent shaping to high-dimensional general-sum games with temporally extended actions and long time horizons. It does so, by simplifying MFOS and effectively capturing both intra-episode and inter-episode information.
@@
Melting Pot 2.0 (Agapiou et al.,, 2023) introduces a comprehensive suite of multi-agent reinforcement learning environments that focus on social interactions and coordination challenges, providing a valuable benchmark for evaluating the scalability and effectiveness of reinforcement learning algorithms in complex, cooperative-competitive settings. The Negotiation Game, introduced by DeVault et al., (2015); Lewis et al., (2017) and subsequently refined by Cao et al., (2018), has proven to be a significant benchmark for studying general-sum games. It integrates elements of strategy and social dilemmas, necessitating that agents balance cooperation and competition to optimize their outcomes. Noukhovitch et al., (2021) analyze this complex benchmark, underscoring its importance in the field. Future investigations will turn towards an even more sophisticated simulation proposed by Zhang et al., (2022), which involves negotiations among countries and regions with diverse resource distributions and preferences in addressing climate change.
@@ 7 Conclusion
In this work, we introduced Advantage Alignment, a novel family of algorithms designed to address the fundamental challenge of achieving self-interested cooperation in multi-agent reinforcement learning, particularly in social dilemmas. By deriving our algorithms from first principles, we distilled opponent shaping to its core components, providing a simple yet powerful mechanism to align agents’ advantages and foster mutually beneficial behaviors. Our approach unifies and generalizes existing opponent shaping methods, such as LOLA and LOQA, demonstrating that they implicitly perform Advantage Alignment through different mechanisms. This unification not only simplifies the mathematical formulation of opponent shaping but also reduces computational complexity, enabling more efficient and scalable algorithms.
Our experiments across a range of social dilemmas, including the Iterated Prisoner’s Dilemma, Coin Game, and a continuous action variant of the Negotiation Game, demonstrate that Advantage Alignment consistently achieves state-of-the-art cooperation and robustness against exploitation. Notably, we extended our methods to complex, large-scale, general-sum environments like Melting Pot’s Commons Harvest Open, addressing challenges that arise from partial observability, high-dimensional observations, and multi-agent interactions. In these settings, Advantage Alignment agents learned sophisticated strategies that balance individual and collective interests, showcasing the potential of our algorithms to scale to real-world applications.
The significance of our work lies in providing a principled, efficient, and scalable solution to the longstanding problem of self-interested cooperation in general-sum games. By enabling agents to autonomously align their interests with one another, Advantage Alignment paves the way for more harmonious and socially beneficial interactions in artificial intelligence systems integrated into human decision-making processes. This has profound implications for the development of AI agents in diverse domains, from autonomous vehicles navigating shared environments to AI assistants collaborating with humans and other agents.
8 Acknowledgements
J.D. would like to thank Michael Noukhovitch for the discussions related to the Negotiation Game. We would also like to acknowledge the technical support of Olexa Bilaniuk. This work was financially supported by A.C.’s CIFAR Canadian AI chair and Canada Research Chair in Learning Representations that Generalize Systematically.